OpenStudy (anonymous):

Evaluate :3^2 + (6 − 2) • 4 − 6/3 =

OpenStudy (anonymous):

is the answer 23

OpenStudy (anonymous):

Have you heard of PEMDAS? P: Parenthesis E: Exponents M: Multiplication D: Division A: Addition S: Subtraction If you have an equation like that, you just go in order: 1) Parenthesis \[3^2+(6-2)*4-\frac{ 6 }{ 3 }\] turns into \[3^2+4*4-\frac{ 6 }{ 3 }\] 2) Exponents \[3^2+4*4-\frac{ 6 }{ 3 }\] turns into \[9+4*4-\frac{ 6 }{ 3 }\] 3) Multiplication \[9+4*4-\frac{ 6 }{ 3 }\] turns into \[9+16-\frac{ 6 }{ 3 }\] 4) Division \[9+16-\frac{ 6 }{ 3 }\] turns into \[9+16-2\] 5) Addition \[9+16-2\] turns into \[25-2\] 6) Subtraction \[25-2\] turns into \[23\] And there you have it! Your answer is correct.

OpenStudy (anonymous):

thanks lol

OpenStudy (harsimran_hs4):

@baracudabrain sometimes or rather mostly text use BODMAS to talk about this concept brackets of division multiplication addition subtraction one and the same thing though

OpenStudy (skullpatrol):

@harsimran_hs4 What is "brackets of division" suppose to mean?

OpenStudy (skullpatrol):

@baracudabrain Don't waste your time memorizing "PEMDAS" or "BODMAS," instead understand what the order of operations are instructing you to do and in particular what order they want you to do it (and then memorize that).

OpenStudy (anonymous):

I do know the order of operation what confused me was the 6/3 I thought it was fraction at first cus the 6 was on top of the 3 not side by side but then I figured it was division

OpenStudy (skullpatrol):

Summary of Order of Operations: 1. First simplify expressions within grouping symbols. 2. Then simplify powers. 3. Then do all multiplications and divisions in order from left to right. 4. Then do all additions and subtractions in order from left to right.

OpenStudy (skullpatrol):